OpenStudy (anonymous):

A 12-foot ladder rests against the side of a house. The base of the ladder is 3 feet away from the side of the house. How high above the ground is the top of the ladder? Round to the nearest tenth of a foot. A. 9.1 feet B. 11.6 feet C. 12.4 feet D. 153.0 feet

OpenStudy (anonymous):

i think A but not sure

Mehek (mehek14):

 \(12^2-3^2=x^2\\144-9=135 \\ \sqrt{135}= 11.6\)

Mehek (mehek14):

you have to use the pythagorean theorem

OpenStudy (anonymous):

thank u

Mehek (mehek14):

yw \(\ddot\smile\)